2. Calculate the AH for the reaction:



CS2 (1) + 2 O2 (g) CO2 (g) + 2 SO2 (g)



Given:



AHCO2 (g) = -393.5 kJ/



mol;



AHSO2 = -296.8 kJ/mol;



AHCS2 (1) = 87.9kJ/mol

1
Expert's answer
2022-03-14T18:15:06-0400

CS2 + 2O2(g) = CO2(g) + 2SO2(g)

"\Delta" Hreaction = "\Delta" HCS2 + "\Delta" HCO2 + 2 * "\Delta" HSO2 = 87.9 + (-393.5) + 2 * (-296.8) = 87.9 - 393.5 - 593.6 = -899.2 kJ/mol.

Answer: -899.2 kJ/mol.